# ReNF: Rethinking the Design Space of Neural Long-Term Time Series Forecasters

### v2

@@ Abstract
Neural Forecasters (NFs) are a cornerstone of Long-term Time Series Forecasting (LTSF). However, progress has been hampered by an overemphasis on architectural complexity at the expense of fundamental forecasting principles. In this work, we return to first principles to redesign the LTSF paradigm. We begin by introducing a Multiple Neural Forecasting Theorem that provides a theoretical basis for our approach. We propose Boosted Direct Output (BDO), a novel forecasting strategy that synergistically combines the advantages of both Auto-Regressive (AR) and Direct Output (DO). In addition, we stabilize the learning process by smoothly tracking the model’s parameters. Extensive experiments show that these principled improvements enable a simple MLP to achieve state-of-the-art performance, outperforming recent, complex models in nearly all cases, without any specific considerations in the area. Finally, we empirically verify our theorem, establishing a dynamic performance bound and identifying promising directions for future research. The code for review is available at: .
@@ 1 Introduction
  • We propose the Multiple Neural Forecasting Theorem (MNFT) with empirical evaluations, which provides a theoretical foundation and rigorously demonstrates the value of employing multiple Neural Forecasters (NFs) in Long-Term Time Series Forecasting (LTSF).
@@ 2 Method
In this section, we lay out the theoretical and methodological foundations of our work. We begin by introducing a core theorem that provides the theoretical underpinnings for our approach. To build intuition, we then illustrate our designs using a simple MLP architecture on carefully selected datasets that highlight the key effects of our methods. The comprehensive experimental setup and full results on benchmark datasets are detailed in Sec.3 and the Appendix.
@@ 2.1 Preliminaries
2.2 Approximate Theorem for LTSF
The following theorem provides the theoretical motivation for our work. For simplicity, we just present the proof for the univariate case; however, the theorem can be readily extended to the multivariate setting.
Figure 1: Illustration of the MNFT.
Theorem 1 (Multiple Neural Forecasting Theorem (MNFT)).
@@ 2.2 Approximate Theorem for LTSF
The proof of the theorem can be found in Appendix C. To interpret, the shown bound value provides us with some intuitions: for instance, the forecasting error of the NFM naturally grows with the forecast horizon $T$ and the data range $\lambda$, because of the expanded solution space.
The most critical insight of the theorem is that an accurate forecast is theoretically achievable even with a rather weak generator, provided a sufficiently large number of candidate outputs is produced. Furthermore, higher-quality historical data and a more powerful NFM (which reduces the estimation error $b$) both serve to constrain the solution space, leading to better predictions. Inspired by this, we can propose the following conception,
@@ 3.2 Ablation Study
By config.1, the BDO reduces to DO with the model structure unchanged. The results, shown in Fig. 7, demonstrate that even on the ETTh1 dataset, which has limited data and is difficult to optimize with deep representations, our BDO strategy can still yield superior performance, especially for very long-term forecasts. This finding empirically supports our claim, derived from Theorem 2.2, that leveraging multiple, hierarchically-generated sub-forecasts provides valuable mutual information that enhances the final prediction.
@@
3.3 Empirical Evaluation of the Theorem
@@
To empirically evaluate our Multiple Neural Forecasting Theorem (MNFT), we implement the post-combination strategy described in Sec. 2.2. This requires defining a combination function, $g_{c}$. While finding an optimal solution without future knowledge is difficult, we can establish a theoretical upper bound on performance. Given the ground truth labels, a trivial yet optimal strategy is to select, at each timestep, the forecast value from among all candidates that has the smallest error. By applying this ”oracle” combination, we aim to explore two points: 1) to quantify the potential accuracy improvement achievable through post-combination, thereby establishing a dynamic empirical bound for our forecasting model; and 2) to verify that this empirical bound behaves in a manner consistent with the intuitions of the MNFT 2.2.
@@
Furthermore, Fig. 8(b) shows how this empirical bound varies with the number of sub-forecasts $K$. The bound consistently decays as $K$ increases, even when the model’s performance saturates. This result aligns perfectly with the core insight from MNFT: the theoretical performance limit improves as the number of candidate forecasts grows.
@@
In a nutshell, our evaluation of the MNFT clarifies two primary roles for Neural Networks in this domain. First, NNs should be leveraged to learn a powerful post-combination function over multiple candidate forecasts. Second, more powerful base forecasters must be developed to better approximate the true data-generating distribution, directly reducing the predictive bias $b$ as identified in our theorem 2.2.
@@
We propose a fundamental theorem to support the mechanism of multiple forecasts. Based on this theorem, we introduce a novel paradigm that combines the advantages of current forecasting methods, enabling more accurate predictions in a hierarchical representation space. To remedy training instability, we leverage EMA to smooth the learning process of neural forecasters. Our empirical results show that with these techniques, a simple MLP can outperform recent SOTA models with significantly less complexity. The value of our theorem is also evaluated and confirmed by experimental results, clearly identifying the role of the Neural Network in building the Long-Term Time Series forecasters.
@@ 5 Reproducibility statement
We have put a lot of effort into ensuring the reproducibility of this work, including the code in the anonymous repository , a complete data description in the Appendix B, and clear explanations of the assumptions and proof of the proposed theorem in the Appendix C.
@@ Appendix A Related Work
We note that ReNF is not the first work to address LTSF using multiple short-term forecasts. A pioneering and highly relevant study by (Nguyen & Chan, 2004) proposed a machine called MNN, which generates long-term forecasts using multiple neural networks, each responsible for a specific interval of the output. The essential distinction between MNN and our approach is that we generate sub-forecasts recursively, with the explicit goal of injecting causality into the DO strategy. In contrast, each sub-network in MNN functions as a standalone AR(1) model. This difference stems from our distinct motivation: whereas MNN was designed primarily to mitigate the error accumulation of one-step-ahead NN models, our work begins with a foundational forecasting theorem. Based on this theorem, we leverage modern deep learning techniques to effectively utilize multiple sub-forecasts, achieving strong empirical results on a wide range of real-world datasets, which constitutes one of our main contributions.
@@
Forecast Combinations. Forecast combination is a classic technique for improving forecast accuracy and robustness by leveraging the diverse strengths of multiple models (Clemen, 1989). In this work, we extend this concept (Bates & Granger, 1969) to the domain of deep learning for long-term time series forecasting. Rather than combining distinct, parallel forecasters into a hybrid model (Zhang, 2003), our framework achieves this goal efficiently within a single neural network. By recursively stacking sub-forecasts, similar in spirit to N-BEATS (Oreshkin et al., 2019), we ensure that the representation learning is both deep and diverse, allowing the LSTF task to benefit from the full power of modern machine learning.
@@ Appendix B Details of Datasets
Appendix C Proof of the approximate theorem
@@ E.3 Further Explorations
Two Factors of BDO. We wish to clarify that our BDO paradigm is comprised of two distinct and essential mechanisms. The first is the addition of a linear head to each block, which generates an explicit sub-forecast from the intermediate representation. The second is the recursive concatenation of this sub-forecast with the input for the subsequent stage, which encourages the network to implicitly learn a post-combination function as motivated by MNFT. To disentangle their respective contributions, we conduct an ablation study on these two factors.
@@ Appendix F Limitation and Further Discussion
The Multiple Neural Forecasting Theorem (MNFT) presented in this paper is foundational but preliminary. While it provides the core intuition for our work, a more rigorous and deeper exploration of its theoretical properties or deriving other related theorems could inspire new and promising research directions.
Specifically, under the paradigm of MNFT, the role of Neural Network (NN) in this area becomes transparent. First and perhaps most importantly, we should leverage NN’s capability to create a powerful post-combination function. While our BDO paradigm is effective, the current recursive strategy for combining sub-forecasts is not fully optimal. This is evident from the performance gap between our final forecast and the theoretical empirical bound as the number of stages increases. Developing more intricate methods to better leverage the full set of sub-forecasts could lead to substantial accuracy gains. Furthermore, the benefit of BDO is less pronounced on certain datasets, such as ETTm2. The underlying reasons for this variance warrant further investigation.
Second, we are consistently supposed to build more powerful Neural Forecasting Machines (NFMs) to approximate the expected distributions of future data, thereby reducing the bias $b$ in the MNFT C. Therefore, a comprehensive study is needed to verify the effects of our proposed techniques when applied to other advanced model architectures beyond MLPs.

### v3

@@ Abstract
Neural Forecasters (NFs) are a cornerstone of Long-term Time Series Forecasting (LTSF). However, progress has been hampered by an overemphasis on architectural complexity at the expense of fundamental forecasting principles. In this work, we return to first principles to redesign the LTSF paradigm. We begin by introducing a Multiple Neural Forecasting Principle that provides a theoretical motivation for our approach. We propose Boosted Direct Output (BDO), a novel forecasting paradigm that synergistically hybridizes the causal nature of Auto-Regressive (AR) models with the stability of Direct Output (DO). In addition, we stabilize the learning process by smoothly tracking the model’s parameters. Extensive experiments show that these principled improvements enable a simple MLP to achieve state-of-the-art performance, outperforming recent, complex models in nearly all cases, without any specific considerations in the area. Finally, we empirically verify our principle, establishing a dynamic performance bound and identifying promising directions for future research. The code for review is available at: .
@@ 1 Introduction
  • We propose the Multiple Neural Forecasting Principle (MNFP) with empirical evaluations, which provides a formal motivation for employing multiple Neural Forecasters (NFs) in Long-Term Time Series Forecasting (LTSF).
@@ 2 Method
In this section, we lay out the theoretical and methodological foundations of our work. We begin by introducing a core proposition that provides the theoretical underpinnings for our approach. To build intuition, we then illustrate our designs using a simple MLP architecture on carefully selected datasets that highlight the key effects of our methods. The comprehensive experimental setup and full results on benchmark datasets are detailed in Sec.3 and the Appendix.
@@ 2.1 Preliminaries
2.2 Theoretical Motivation for Multiple Forecasts
The following proposition provides the theoretical motivation for our work. For simplicity, we just present the proof for the univariate case; however, the proposition can be readily extended to the multivariate setting.
Figure 1: Illustration of the MNFP.
Proposition 1 (Multiple Neural Forecasting Proposition (MNFP)).
@@ 2.2 Theoretical Motivation for Multiple Forecasts
The proof can be found in Appendix C. To interpret, the shown bound value provides us with some intuitions: for instance, the forecasting error of the NFM naturally grows with the forecast horizon $T$ and the data range $\lambda$, because of the expanded solution space.
The most critical insight of the proposition is that an accurate forecast is theoretically achievable even with a rather weak generator, provided a sufficiently large number of candidate outputs is produced. This principle is rooted in the well-established theory of ensemble methods (Bates & Granger, 1969), which demonstrates that combining multiple diverse estimators can lead to a more accurate and robust final estimate. Furthermore, higher-quality historical data and a more powerful NFM (which reduces the estimation error $b$) both serve to constrain the solution space, leading to better predictions.
@@ 3.2 Ablation Study
By config.1, the BDO reduces to DO with the model structure unchanged. The results, shown in Fig. 7, demonstrate that even on the ETTh1 dataset, which has limited data and is difficult to optimize with deep representations, our BDO strategy can still yield superior performance, especially for very long-term forecasts. This finding empirically supports our claim, derived from Proposition 2.2, that leveraging multiple, hierarchically-generated sub-forecasts provides valuable mutual information that enhances the final prediction.
@@
3.3 Empirical Evaluation of the Proposition
@@
To empirically evaluate our Multiple Neural Forecasting Proposition, we implement the post-combination strategy described in Sec. 2.2. This requires defining a combination function, $g_{c}$. While finding an optimal solution without future knowledge is difficult, we can establish a theoretical upper bound on performance. Given the ground truth labels, a trivial yet optimal strategy is to select, at each timestep, the forecast value from among all candidates that has the smallest error. By applying this ”oracle” combination, we aim to explore two points: 1) to quantify the potential accuracy improvement achievable through post-combination, thereby establishing a dynamic empirical bound for our forecasting model; and 2) to verify that this empirical bound behaves in a manner consistent with the intuitions of the MNFP 2.2.
@@
Furthermore, Fig. 8(b) shows how this empirical bound varies with the number of sub-forecasts $K$. The bound consistently decays as $K$ increases, even when the model’s performance saturates. This result aligns perfectly with the core insight from MNFP: the theoretical performance limit improves as the number of candidate forecasts grows.
@@
In a nutshell, our evaluation of the MNFP clarifies two primary roles for Neural Networks in this domain. First, NNs should be leveraged to learn a powerful post-combination function over multiple candidate forecasts. Second, more powerful base forecasters must be developed to better approximate the true data-generating distribution, directly reducing the predictive bias $b$ as identified in our proposition 2.2.
@@
We propose a fundamental proposition to support the mechanism of multiple forecasts. Based on this proposition, we introduce a novel paradigm that combines the advantages of current forecasting methods, enabling more accurate predictions in a hierarchical representation space. To remedy training instability, we leverage EMA to smooth the learning process of neural forecasters. Our empirical results show that with these techniques, a simple MLP can outperform recent SOTA models with significantly less complexity. The value of our proposition is also evaluated and confirmed by experimental results, clearly identifying the role of the Neural Network in building the Long-Term Time Series forecasters.
@@ 5 Reproducibility statement
We have put a lot of effort into ensuring the reproducibility of this work, including the code in the anonymous repository , a complete data description in the Appendix B, and clear explanations of the assumptions and proof of the proposition in the Appendix C.
@@ Appendix A Related Work
We note that ReNF is not the first work to address LTSF using multiple short-term forecasts. A pioneering and highly relevant study by (Nguyen & Chan, 2004) proposed a machine called MNN, which generates long-term forecasts using multiple neural networks, each responsible for a specific interval of the output. The essential distinction between MNN and our approach is that we generate sub-forecasts recursively, with the explicit goal of injecting causality into the DO strategy. In contrast, each sub-network in MNN functions as a standalone AR(1) model. This difference stems from our distinct motivation: whereas MNN was designed primarily to mitigate the error accumulation of one-step-ahead NN models, our work begins with a foundational forecasting proposition. Based on this proposition, we leverage modern deep learning techniques to effectively utilize multiple sub-forecasts, achieving strong empirical results on a wide range of real-world datasets, which constitutes one of our main contributions.
@@
Forecast Combinations. Forecast combination is a classic technique for improving forecast accuracy and robustness by leveraging the diverse strengths of multiple models (Clemen, 1989). In this work, we extend this concept (Bates & Granger, 1969) to the domain of deep learning for long-term time series forecasting. Rather than combining distinct, parallel forecasters into a hybrid model (Zhang, 2003), our framework achieves this goal efficiently within a single, structured approach for generating and implicitly combining forecasts within a single neural network. By recursively stacking sub-forecasts, similar in spirit to N-BEATS (Oreshkin et al., 2019), we ensure that the representation learning is both deep and diverse, allowing the LSTF task to benefit from the full power of modern machine learning.
@@ Appendix B Details of Datasets
Appendix C Proof of the approximate proposition
@@ E.3 Further Explorations
Two Factors of BDO. We wish to clarify that our BDO paradigm is comprised of two distinct and essential mechanisms. The first is the addition of a linear head to each block, which generates an explicit sub-forecast from the intermediate representation. The second is the recursive concatenation of this sub-forecast with the input for the subsequent stage, which encourages the network to implicitly learn a post-combination function as motivated by MNFP. To disentangle their respective contributions, we conduct an ablation study on these two factors.
@@ Appendix F Limitation and Further Discussion
The Multiple Neural Forecasting Proposition (MNFP) presented in this paper is foundational but preliminary. While it provides the core intuition for our work, a more rigorous and deeper exploration of its theoretical properties or deriving other related theorems could inspire new and promising research directions.
Specifically, under the paradigm of MNFP, the role of Neural Network (NN) in this area becomes transparent. First and perhaps most importantly, we should leverage NN’s capability to create a powerful post-combination function. While our BDO paradigm is effective, the current recursive strategy for combining sub-forecasts is not fully optimal. This is evident from the performance gap between our final forecast and the theoretical empirical bound as the number of stages increases. Developing more intricate methods to better leverage the full set of sub-forecasts could lead to substantial accuracy gains. Furthermore, the benefit of BDO is less pronounced on certain datasets, such as ETTm2. The underlying reasons for this variance warrant further investigation.
Second, we are consistently supposed to build more powerful Neural Forecasting Machines (NFMs) to approximate the expected distributions of future data, thereby reducing the bias $b$ in the MNFP C. Therefore, a comprehensive study is needed to verify the effects of our proposed techniques when applied to other advanced model architectures beyond MLPs.